Design a phase shift oscillator to oscillate at 2 kHz using an OPAMP. Given C=0.1



µF.




1
Expert's answer
2022-02-10T09:42:09-0500

A 3-stage RC shift oscillator is required to produce an oscillation frequency of 2 kHz


"f=\frac{1}{2\pi\>RC\sqrt{2N}}"


"2=\frac{1}{2\pi\>R(0.1×10^{-6})\sqrt{2×3}}"


"R=\frac{1}{2\pi×2×10^3(0.1×10^{-6})\sqrt{6}}"

"=324.9\Omega"


The resistor in each stage should have a resistance of "324.9\Omega"

And a capacitor of "0.1\mu" F